OpenStudy (anonymous):

Solve by the substitution method. 5x+9y=10 x=18-5y. What is the solution as an odered pair?

OpenStudy (anonymous):

5(18-5y)+9y=10 y=-80/-16=5 x=18-5*5=-7 so (-7,5) should be your answer if my arithmetic is correct